OpenStudy (anonymous):

Solve: −2(x − 5) = −4

OpenStudy (anonymous):

Divide by -2 both sides

OpenStudy (anonymous):

x - 5 = 2 x = 7

OpenStudy (compassionate):

-2x + 10 = - 4 -2x = - 14 --- --- -2 -2 x = 7

OpenStudy (compassionate):

@Yahoo! return the favor c: